OpenStudy (anonymous):

@Nnesha - Can you Help?: What is the exact value of tan 75 degrees? Not just answers please, explain how to do it... PLEASE HELP!!!

ganeshie8 (ganeshie8):

Hint : \(\large 75 = 30+45\)

ganeshie8 (ganeshie8):

look up \(\tan(A+B)\) formula in ur notes

OpenStudy (anonymous):

ok so with the half angle formula right?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I get this: tan 75 = tan (150)/2 = 1 - cos(theta)/sin (theta) = 1-(-sqrt(3)/2)/(1/2) = 2 + sqrt (3)

OpenStudy (anonymous):

Looks a little weird but I think its right??

Nnesha (nnesha):

i would use this one \[\huge\rm tan(\color{reD}{a}+\color{blue}{b})=\frac{ \tan \color{red}{a}+ \tan \color{blue}{b} }{ 1-tan\color{reD}{a}tan\color{blue}{b} }\] sum formula for tan

Nnesha (nnesha):

a ,b are two numbers on the unit circle that add up to 75 degree

Nnesha (nnesha):

look above 't first comment by ganeshie8 30+45=75 so a=30 b=45

OpenStudy (anonymous):

got it thanks!

Nnesha (nnesha):

np